# Hook Length Biases in $t$-Core Partitions

### 10 Mar 2026 v1

@@ Abstract.
Recently, the theory of hook length biases has emerged as a prominent research topic. Led by Ballantine, Burson, Craig, Folsom, and Wen [Res. Math. Sci., 2023], hook length biases are being explored for ordinary partitions, odd versus distinct partitions, self-conjugate versus distinct odd partitions. Lately, Singh and Barman [J. Number Theory, 2024] opened the door to hook length biases in $\ell$-regular partitions. In this work, we extend the theory of hook length biases to $t$-core partitions. For example, let $a_{t,k}(n)$ denote the number of hooks of length $k$ in all $t$-core partitions of $n$, then we find that $a_{3,1}(n)\geq a_{3,2}(n)\geq a_{3,4}(n)$ and $a_{4,1}(n)\geq a_{4,3}(n)$ for all $n$. The methods employed in this work are combinatorial.
@@ 1. Introduction
One of the primary motivations for studying $t$-core partitions is their deep connection to the representation theory of symmetric groups. Specifically, $t$-core partitions determine the block structure of the representations of symmetric groups. For further aspects of $t$-core partitions in representation theory of symmetric groups, see [7, 9, 12]. Furthermore, $t$-core partitions also have a significant relation to the quadratic forms. This relationship is more fascinating when the above generating function is a weight-3/2 modular form. For example, Ono and Sze [13] proved that if $8n+5$ is square-free then $c_{4}(n)=1/2h(-32n-20)$, where $h(D)$ denotes the discriminant $D$ class number, viz. the order of the class group of the discriminant $D$ binary quadratic forms.
@@
Recently, inspired in part by hook-content formulas for specific restricted partitions in representation theory, Ballantine et al. [1] investigated the total number of hooks of a given length in odd versus distinct partitions. Their work further motivated Singh and Barman [16] to study hook length biases in ordinary and regular partitions. For integers $t\geq 2$ and $k\geq 1$, let $b_{t,k}(n)$ denote the number of hooks of length $k$ in all the $t$-regular partitions of $n$. Singh and Barman found that $b_{2,2}(n)\geq b_{2,1}(n)$ for all $n>4$, while $b_{2,2}(n)\geq b_{2,3}(n)$ for all $n\geq 0$. They also conjectured that $b_{3,2}(n)\geq b_{3,1}(n)$ for all $n\geq 28$, which was first settled by He and Liu [8]. He and Liu also proved that $b_{2,4}(n)\geq b_{2,3}(n)$ for all $n\geq 82$. Investigating another query by Singh and Barman, Qu and Zang [14] found that $b_{2,4}(n)\geq b_{2,5}(n)$ for all $n\geq 0$ and $n\neq 5$, and $b_{2,6}(n)\geq b_{2,7}(n)$ when $n\neq 7$.
@@
The following conjecture arises naturally.
Conjecture 1.9.
@@
In fact, we know that $a_{2,1}(n)=\ell$, $a_{2,3}(n)=\ell-1$ for $n=\ell(\ell+1)/2$. To prove Conjecture 1.9, it therefore suffices to show that $a_{4,3}(n)\geq\ell$ for $n=\ell(\ell+1)/2$. The form A in Figure 11 in Section 4 implies $a_{4,3}(n)\geq\ell-1$. Consequently, the task reduces to proving the existence of at least two 4-core partitions of $n$, whenever $n=\ell(\ell+1)/2$. The following alternative, yet elementary, approach may also settle the problem. Note that
Due to the widely popular Product-to-Sum identity:
$\prod_{j=1}^{\infty}\dfrac{(1-q^{2j})^{2}}{1-q^{j}}=\sum_{m=0}^{\infty}q^{m(m+1)/2},$
we have
$=\sum_{m=0}^{\infty}q^{m(m+1)/2}+\sum_{\begin{subarray}{c}m,r,s=0\\ r=s\neq 0\end{subarray}}^{\infty}q^{m(m+1)/2+r(r+1)+s(s+1)}.$
In the above identity, we already see exponents of the form $m(m+1)/2$ in the first summation. Therefore, the objective reduces to confirming that for every $h\geq 2$, there are integers $m,r,s$ such that
$\dfrac{h(h+1)}{2}$ $=\dfrac{m(m+1)}{2}+r(r+1)+s(s+1),$
equivalently,
$(2h+1)^{2}+4$ $=(2m+1)^{2}+2(2r+1)^{2}+2(2s+1)^{2}.$

### 12 Mar 2026 v2

@@ Abstract.
Recently, the theory of hook length biases has emerged as a prominent research topic. Led by Ballantine, Burson, Craig, Folsom, and Wen [Res. Math. Sci., 2023], hook length biases are being explored for ordinary partitions, odd versus distinct partitions, self-conjugate versus distinct odd partitions. Lately, Singh and Barman [J. Number Theory, 2024] opened the door to hook length biases in $\ell$-regular partitions. In this work, we extend the theory of hook length biases to $t$-core partitions. For example, let $a_{t,k}(n)$ denote the number of hooks of length $k$ in all $t$-core partitions of $n$, then we find that $a_{3,1}(n)\geq a_{3,2}(n)\geq a_{3,4}(n)$ and $a_{4,1}(n)\geq a_{4,3}(n)$ for all $n$. The methods employed in this work are mainly combinatorial.
@@ 1. Introduction
One of the primary motivations for studying $t$-core partitions is their deep connection to the representation theory of symmetric groups. Specifically, $t$-core partitions determine the block structure of the representations of symmetric groups. For further aspects of $t$-core partitions in representation theory of symmetric groups, see [8, 10, 13]. Furthermore, $t$-core partitions also have a significant relation to the quadratic forms. This relationship is more fascinating when the above generating function is a weight-3/2 modular form. For example, Ono and Sze [14] proved that if $8n+5$ is square-free then $a_{4}(n)=1/2h(-32n-20)$, where $h(D)$ denotes the discriminant $D$ class number, viz. the order of the class group of the discriminant $D$ binary quadratic forms.
@@
Recently, inspired in part by hook-content formulas for specific restricted partitions in representation theory, Ballantine, Burson, Craig, Folsom, and Wen [1] investigated the total number of hooks of a given length in odd versus distinct partitions. Their work further motivated Singh and Barman [17] to study hook length biases in ordinary and regular partitions. For integers $t\geq 2$ and $k\geq 1$, let $b_{t,k}(n)$ denote the number of hooks of length $k$ in all the $t$-regular partitions of $n$. Singh and Barman found that $b_{2,2}(n)\geq b_{2,1}(n)$ for all $n>4$, while $b_{2,2}(n)\geq b_{2,3}(n)$ for all $n\geq 0$. They also conjectured that $b_{3,2}(n)\geq b_{3,1}(n)$ for all $n\geq 28$, which was first settled by He and Liu [9]. He and Liu also proved that $b_{2,4}(n)\geq b_{2,3}(n)$ for all $n\geq 82$. Investigating another query by Singh and Barman, Qu and Zang [15] found that $b_{2,4}(n)\geq b_{2,5}(n)$ for all $n\geq 0$ and $n\neq 5$, and $b_{2,6}(n)\geq b_{2,7}(n)$ when $n\neq 7$.
@@
The following theorem arises as a natural extension, which presents biases among fixed hooks in different core partitions.
Theorem 1.9.
@@
In the following sections, we prove our results.
@@
5. Proof of Theorem 1.9
From Proposition 1.1, we know that $a_{2,1}(n)=\ell$, $a_{2,3}(n)=\ell-1$ for $n=\ell(\ell+1)/2$. To prove Theorem 1.9, it therefore suffices to show that $a_{4,3}(n)\geq\ell$ for $n=\ell(\ell+1)/2$. The form A in Figure 11 in Section 4 implies $a_{4,3}(n)\geq\ell-1$. Consequently, the task reduces to proving the existence of at least two 4-core partitions of $n$, whenever $n=\ell(\ell+1)/2$. The following alternative, yet elementary, approach settles the problem. Note that
Due to the second equality in (2.1), we have
$=\sum_{m=0}^{\infty}q^{m(m+1)/2}+\sum_{\begin{subarray}{c}m,r,s=0\\ \textup{$r,s\neq 0$ simultaneously}\end{subarray}}^{\infty}q^{m(m+1)/2+r(r+1)+s(s+1)}.$
In the above identity, we see exponents of the form $m(m+1)/2$ in the first summation. This means that the first summation contributes one 4-core partition for every triangular number. Therefore, the objective reduces to confirming that for every $h\geq 2$, there are integers $m,r,s$ such that
$\dfrac{h(h+1)}{2}$ $=\dfrac{m(m+1)}{2}+r(r+1)+s(s+1),$
equivalently,
$(2h+1)^{2}+4$ $=(2m+1)^{2}+2(2r+1)^{2}+2(2s+1)^{2}.$
Dickson [6, Theorem VI] proved that the quadratic form $x^{2}+2y^{2}+2z^{2}$ represents all positive integers not of the form $4^{k}(8n+7)$.
Now, for any $h\geq 2$, we have
$(2h+1)^{2}+4\equiv 5\pmod{8}$,
which readily implies that $(2h+1)^{2}+4$ is not of the form $4^{k}(8n+7)$, as $4^{k}(8n+7)\equiv 0\textup{~or~}4\pmod{8}$. Therefore, by Dickson’s result [6, Theorem VI], there exist integers $x,y,z$ such that $(2h+1)^{2}+4=x^{2}+2y^{2}+2z^{2}$. It remains to show that $x,y$, and $z$ are odd. From the above arguments, we have
(5.1) $x^{2}+2y^{2}+2z^{2}\equiv 5\pmod{8}$.
Since $2y^{2}+2z^{2}$ is even, $x^{2}$ must be odd, which implies that $x$ is odd. Thus, equation (5.1) reduces to
$y^{2}+z^{2}\equiv 2\pmod{4}$.
Now it is routine that a case by case analysis of the parities of $y$ and $z$ implies that the only way the above holds is when $y$ and $z$ are both odd. Thus, we have proved that for every $h\geq 2$ there exist integers $m,r,s$ such that
$(2h+1)^{2}+4=(2m+1)^{2}+2(2r+1)^{2}+2(2s+1)^{2}.$
This completes the proof of the theorem. ∎